Suppose that business travelers and tourists have the following demand for airline

tickets from Lahore to Karachi.

price Quantity demanded

(business travelers)


Quantity demanded

(tourists)

$150 2100 1000

200 2000 800

250 1900 600

300 1800 400

As the price of ticket rises from $200 to $300, what is the price elasticity of demand for


(i) Business travelers

(ii) Tourists

(Use the midpoint method in your calculations.)


1
Expert's answer
2021-01-29T20:00:15-0500

(i) Let's calculate the price elasticity of demand for business travelers:


"E_d=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},""E_d=\dfrac{\dfrac{1800-2000}{0.5(1800+2000)}}{\dfrac{300-200}{0.5(300+200)}}=-0.26"

(ii) Let's calculate the price elasticity of demand for tourists:


"E_d=\dfrac{\dfrac{Q_2-Q_1}{0.5(Q_2+Q_1)}}{\dfrac{P_2-P_1}{0.5(P_2+P_1)}},""E_d=\dfrac{\dfrac{400-800}{0.5(800+400)}}{\dfrac{300-200}{0.5(300+200)}}=-1.66"